# On the irreducible character degrees of symmetric groups and their multiplicities

### v1

@@ 1 Introduction
The multisets of degrees of the irreducible characters of the symmetric groups $S_{n}$ have been studied for over a century. Over the last 50 years or so there has been work on two particular aspects of this: the largest degrees of irreducible characters, and the multiplicities of irreducible character degrees. Here we will propose a few new conjectures and avenues of research, on the basis of experiments with symmetric groups of degree up to around 120, along with some new results in this area.
We start with multiplicities. If $G$ is a finite group, write $m(G)$ for the maximal multiplicity of an irreducible character degree of $G$, and write $m(n)=m(S_{n})$. In [3] it was proved that $m(n)\to\infty$ as $n\to\infty$, and together with [17] this proves that $m(G)\to\infty$ as $|G|\to\infty$. A natural question is to ask for the growth rate of $m(G)$ in general, and $m(n)$ in particular. A lower bound for the growth of $m(n)$, approximately $n^{1/7}$, was proved in [4, Theorem E], but this is far from optimal, and is a remnant of the methods of [3]. The theorem we prove here is the following, which has been used in [15] to find all finite groups $G$ for which $m(G)=2$. This result is a slight specialization of Theorem 3.2 below.
@@
Since the degrees of the unipotent characters of the general linear groups are related to those of symmetric groups, it is reasonable to suspect that one might be able to say something about multiplicities of character degrees for general linear groups. As we see in Section 2.2, the degrees of the characters associated to a partition and its conjugate, while equal for $S_{n}$, are generally not for $\operatorname{GL}_{n}(q)$, so it is not even clear that there are two unipotent characters with the same degree for $\operatorname{GL}_{n}(q)$. To the author’s knowledge there is no established theory about this in the literature, so this paper starts the investigation of this subject, with the following contribution.
@@
Connected to the maximal multiplicity of a character degree we have the average multiplicity of a character degree. Write $k(G)$ for the number of conjugacy classes of a finite group $G$, or equivalently the number of irreducible characters of $G$. Let $\mathrm{cd}(G)$ denote the set of irreducible character degrees of $G$, and $\mathrm{cc}(G)$ denote the set of conjugacy class sizes of $G$. The average character degree multiplicity is $k(G)/|\mathrm{cd}(G)|$, and the average conjugacy class size multiplicity is $k(G)/|\mathrm{cc}(G)|$. We compute both of these for symmetric groups; the former is calculated up to $S_{115}$, with a pattern emerging (see Figure 3.1) but the asymptotics remaining unclear at this stage. For conjugacy classes, only data up to $S_{85}$ was necessary for a convincing conjecture to emerge (Figure 3.2). The following two conjectures are considered in more detail in Section 3, and the concepts have been considered before, for example in [10, 17].
@@
There exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, if $G$ is a finite group with no alternating composition factor, then $|G|\leq f(k(G)/|\mathrm{cd}(G)|)$.
@@
There exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, if $G$ is a finite group then $|G|\leq f(k(G)/|\mathrm{cc}(G)|)$.
@@
At the end of Section 4 we also include some information about how far the partitions $\lambda^{(n)}$ witnessing $b_{1}(n)$ stray from being self-conjugate. Contrary to what might be considered reasonable – that as $n$ increases the partition becomes closer and closer to self-conjugate (mimicking the curve known to maximize the continuous version of the hook length formula found in [14, 21]) – it seems that the difference between $\lambda^{(n)}$ and a self-conjugate partition is unbounded as $n$ grows. This has implications for any posited algorithms for guessing the largest-degree character that start by looking at self-conjugate partitions.
@@ 2.1 Symmetric groups
Notice that a partition and its conjugate have the same multiset of hook lengths.
@@
$\chi_{\lambda}(1)=\frac{n!}{\prod_{(i,j)\in\lambda}h_{i,j}},$
where the product runs over all boxes $(i,j)$ in $\lambda$ and $h_{i,j}$ is the hook length at the box $(i,j)$. Let $H(\lambda)$ denote the multiset of hook lengths of $\lambda$.
At one point we will use the branching rule. If $\lambda$ is a partition of $n$, then a removable node is a box in the Young diagram of $\lambda$ that, upon removal, yields a Young diagram of a (smaller) partition. This is equivalent to the box having hook length $1$. The opposite is an addable node: an addable node is a node that is not in the Young diagram of $\lambda$ but, if it is added to $\lambda$, the result is also the Young diagram of a partition.
The branching rule states that if $\lambda$ is a partition of $n$, then the restriction of $\chi_{\lambda}$ to $S_{n-1}$ is the sum of all irreducible characters $\chi_{\mu}$ such that $\mu$ is obtained from $\lambda$ by removing a removable node. By Frobenius reciprocity, the induction of $\chi_{\lambda}$ to $S_{n+1}$ is the sum of $\chi_{\mu}$ for $\mu$ obtained from $\lambda$ by adding an addable node.
@@ 2.2 General linear groups
The complex character theory of the general linear groups $\operatorname{GL}_{n}(q)$ uses many of the combinatorial concepts involved in symmetric groups such an hook lengths. One may find a description of some of this theory in [2].
We consider a subset of the ordinary characters, called unipotent characters. These are labelled by partitions of $n$, and their degrees are polynomials in $q$. If $\lambda=(\lambda_{1},\dots,\lambda_{r})$ is a partition of $n$, define Lusztig’s $a$-function by
@@
The degree of the character $\chi_{\lambda}$ is
@@
Generic degrees can be found using cyclotomic Hecke algebras, and in the best-case scenario there is a single unipotent character with that generic character degree. Coincidences between generic degrees of unipotent characters make it more difficult to understand the bijection between irreducible unipotent characters in a block of $\operatorname{GL}_{n}(q)$ and the simple modules for the Brauer correspondent given by the cyclotomic Hecke algebra. This is an important part of various formulations of Broué’s conjecture for groups of Lie type, so this question has ‘real-world’ applications [5, 6].
@@ 3.1 Symmetric groups
For (ii), we find $9$ periodic clusters of period $9$, order $8$ and size various $n$, one for each congruence modulo $9$.
@@
It is not known whether there are clusters of order $8$ of these sizes, but it is certainly likely that there are clusters of some of the larger sizes in the list above. There is no cluster of order $8$ and size up to $83$ by an exhaustive computer search, but computer programs take inconveniently long times to prove this (the case $n=82$ took around two weeks using a Magma program [^1] This is why Table 4.1 stops at $n=115$.) because keeping track of all of the multisets of hook lengths is challenging.
@@
Moving from the multisets having to coincide to coincidences of character degrees, we can prove much better bounds than those in Theorem 3.1. First, for $n\leq 115$ we can use Magma to explicitly calculate $m(n)$. It is given in Table 4.1. Combining this with Theorem 3.1 proves almost all of the next result.
@@
For the bound for $m(n)\geq 4$, we can use Theorem 3.1 to prove the result for $n\geq 22$, and then Table 4.1 for the remaining values of $n$. For the second and third parts, Theorem 3.1 proves the result for $n\geq 206$, and from the remark afterwards, we know the result for $n\geq 182$. Table 4.1 gives us the result for $n\leq 115$. At this point one can either check each value in the range $116\leq n\leq 181$ individually that $m(n)\geq 8$ just by calculating all character degrees one at a time until at least eight of the same are found, or use the other known clusters of order $8$ to reduce the number of possible $n$ for which a brute-force search is needed. Either way, this proves the result.
@@
From Table 4.1 we see what the answer should be for larger bounds: for $m(n)\geq 10$ we need $n\geq 25$, for $m(n)\geq 12$ we need $n\geq 28$, and so on. However, this is conjectural because we do not have a version of Theorem 3.1 for these values, to eliminate all but a small number of $n$ that we can check individually. The proof in [3] gives no explicit bounds, and the version in [4] gives us bounds of the form $m(n)\geq O(n^{1/7})$. These are obviously not useful for larger values of $m(n)$.
@@
However, we know that there are coincidences of character degrees, since $m(n)\to\infty$ as $n\to\infty$. Figure 3.1 gives the ratio $k(S_{n})/|\mathrm{cd}(S_{n})|$ for $1\leq n\leq 115$. It looks far from conclusive at this stage whether the average multiplicity has a limit (maybe 2.5?) or tends to infinity as $n$ increases. [^2] In [17], Moretó calculated this ratio for $n\leq 55$, but the general trend is not particularly obvious for those smaller values of $n$.
@@
We can ask the question for arbitrary finite groups, a significant strengthening of the result by Moretó [17] and the author [3] on $|G|$ being bounded by a function of $m(G)$. In [17, p.245], Moretó asks this question specifically for soluble groups rather than all groups, but it looks as though this might be true, at least for all groups other than symmetric groups.
@@
There exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, if $G$ is a finite group with no alternating composition factor, then $|G|\leq f(k(G)/|\mathrm{cd}(G)|)$.
The only reason that alternating groups are excluded from this is that Figure 3.1 does not compel the author to believe that the average multiplicity is unbounded for symmetric groups.
@@
There exists a function $f:\mathbb{N}\to\mathbb{N}$ such that, if $G$ is a finite group then $|G|\leq f(k(G)/|\mathrm{cc}(G)|)$.
In other words, the average multiplicity of a conjugacy class size should tend to infinity as the group order does. The evidence for this for symmetric groups is much more convincing, and looks potentially easier to prove since centralizer sizes are easier to compute than character degrees. The graph, Figure 3.2, is a smooth exponential-like curve as well, suggesting an asymptotically accurate formula might even be findable. [^3] Although it looks like an exponential curve, it cannot of course be strictly exponential because $k(S_{n})$ is not even exponential in $n$.
@@
We note that Conjectures 3.3 and 3.4 hold for $p$-groups and soluble groups with trivial Frattini subgroup [18, Lemmas 2.2 and 2.3], and Lie-type groups of bounded rank by [17, Lemma 6.3]. Conjecture 3.3 is true for simple groups of Lie type by [13, Corollary 1.6].
Having discussed the first column of the character table, one could ask the question about more than one column. A question of Alexander Ryba was whether there are irreducible characters $\lambda$ and $\mu$ of $S_{n}$ for $n\geq 2$ such that $\lambda(1)=\mu(1)$ and $\lambda((1,2))=\mu((1,2))\neq 0$ (so that $\lambda$ and $\mu$ are not conjugate)? The first such case is for $S_{12}$, where the characters associated to the partitions $(7,1,1,1,1,1)$ and $(4,4,4)$ have the same character degree and value on an involution. By $S_{30}$ there are 70 pairs of such partitions, and by $S_{40}$ there are 314. The first $n$ for which there are two characters with the same value on $1$, $(1,2)$ and $(1,2,3,4)$ is $35$.
@@
Let $X$ be a finite set of permutations, and let $k\geq 1$ be an integer. For all sufficiently large $n$, there exist $k$ distinct irreducible characters of $S_{n}$ whose values on permutations $g\in S_{n}$ coincide for all $g$ of cycle type in $X$.

### v3

@@ Abstract
Keywords: representations of finite groups, symmetric groups, character degrees.
@@ 1 Introduction
The multisets of degrees of the irreducible characters of the symmetric groups $S_{n}$ have been studied for over a century. Over the last 50 years or so there has been work on two particular aspects of this: the largest degrees of irreducible characters, and the multiplicities of irreducible character degrees. Here we will propose a few new conjectures and avenues of research, on the basis of experiments with symmetric groups of degree up to either 129 or 150 (depending on the question), along with some new results in this area.
We start with multiplicities. If $G$ is a finite group, write $m(G)$ for the maximal multiplicity of an irreducible character degree of $G$, and write $m(n)=m(S_{n})$. In [3] it was proved that $m(n)\to\infty$ as $n\to\infty$, and together with [20] and the Classification of Finite Simple Groups, this proves that $m(G)\to\infty$ as $|G|\to\infty$, i.e., $|G|$ is bounded above by a function of $m(G)$. A natural question is to ask for the growth rate of $m(G)$ in general, and $m(n)$ in particular. For groups of Lie type $G(q)$ there are fairly easy lower bounds, for example roughly $q/4$ for $G(q)=\operatorname{PSL}_{2}(q)$.
An explicit lower bound for the growth of $m(n)$, approximately $n^{1/7}$, was proved in [4, Theorem E], but this is far from optimal, and the precise nature of this bound is a consequence of the methods of [3]. The theorem we prove here is the following, which has been used in [17] to find all finite groups $G$ for which $m(G)=2$ (the Baby Monster is the largest such group). This result is a slight specialization of Theorem 3.2 below.
@@
The irreducible characters of symmetric groups are described combinatorially, see Section 2.1 below for more information and background. The unipotent characters of the general linear groups, a specific, important subset of the irreducible characters of $\operatorname{GL}_{n}(q)$ that are defined independently of $q$, are parametrized by partitions, as for the irreducible characters of symmetric groups. Their degrees are combinatorially defined in a similar manner to the hook formula for symmetric groups (see Section 2.2 below or [2, p.465] for an alternative version). It is reasonable to suspect that one might be able to say something about multiplicities of character degrees for general linear groups. As we see in Section 2.2, the degrees of the characters associated to a partition and its conjugate, while equal for $S_{n}$, are generally not for $\operatorname{GL}_{n}(q)$, so it is not even clear that there are two unipotent characters with the same degree for $\operatorname{GL}_{n}(q)$. To the author’s knowledge there is no established theory about this in the literature, so this paper starts the investigation of this subject, with the following contribution.
@@
Connected to the maximal multiplicity of a character degree we have the average multiplicity of a character degree. Write $k(G)$ for the number of conjugacy classes of a finite group $G$, or equivalently the number of irreducible characters of $G$. Let $\mathrm{cd}(G)$ denote the set of irreducible character degrees of $G$, and $\mathrm{cc}(G)$ denote the set of conjugacy class sizes of $G$, the ‘dual notion’. The average character degree multiplicity is $k(G)/|\mathrm{cd}(G)|$, and the average conjugacy class size multiplicity is $k(G)/|\mathrm{cc}(G)|$. We compute both of these for symmetric groups $S_{n}$ up to $n=129$; a pattern emerges for $k(G)/|\mathrm{cd}(G)|$ (see Figure 3.1) but the asymptotics remain unclear at this stage. For conjugacy classes, however, data up to $S_{85}$ is already enough for a convincing conjecture to emerge (see Figure 3.2). The following two conjectures are considered in more detail, with current progress on them, in Section 3, and the concepts have been considered before, for example in [12, 20].
@@
There exists a function $f:\mathbb{Q}\to\mathbb{Q}$ such that, if $G$ is a finite group with no alternating composition factor, then $|G|\leq f(k(G)/|\mathrm{cd}(G)|)$.
@@
There exists a function $f:\mathbb{Q}\to\mathbb{Q}$ such that, if $G$ is a finite group then $|G|\leq f(k(G)/|\mathrm{cc}(G)|)$.
@@
Conjecture 4.2.
For any $m\geq 2$, the ratio $b_{m}(n)/b_{1}(n)\to 1$ as $n\to\infty$.
This should be a consequence of there being many ‘large-degree’ characters clustered around the maximal degree, but as far as the author is aware this result is not known.
At the end of Section 4 we also include some information about how far the partitions $\lambda^{(n)}$ witnessing $b_{1}(n)$ stray from being self-conjugate. Contrary to what might be considered at first glance reasonable – that as $n$ increases the partition becomes closer and closer to self-conjugate (mimicking the curve known to maximize the continuous version of the hook length formula found in [16, 24]) – it seems that the difference between $\lambda^{(n)}$ and a self-conjugate partition is unbounded as $n$ grows. This has implications for the posited algorithms for guessing the largest-degree irreducible character that start by looking at self-conjugate partitions, or the largest-degree irreducible character for $S_{n-1}$.
@@
I would like to thank Benjamin Sambale for noting an error in the formulation of Conjectures 3.3 and 3.4. I would like to thank the two referees for their helpful and incisive comments, that improved a number of facets of this work.
@@ 2.1 Symmetric groups
Notice that a partition and its conjugate have the same multiset of hook lengths. We denote this multiset of hook lengths by $H(\lambda)$, and wish to emphasise that it is a multiset, not a set.
@@
$\chi_{\lambda}(1)=\frac{n!}{\prod_{h\in H(\lambda)}h},$
where the product runs over hook lengths in $\lambda$.
At one point we will use the branching rule. If $\lambda$ is a partition of $n$, then a removable box is a box in the Young diagram of $\lambda$ that, upon removal, yields a Young diagram of a (smaller) partition. This is equivalent to the box having hook length $1$. The opposite is an addable box: an addable box is a box that is not in the Young diagram of $\lambda$ but, if it is added to $\lambda$, the result is also the Young diagram of a partition.
The branching rule states that if $\lambda$ is a partition of $n$, then the restriction of $\chi_{\lambda}$ to $S_{n-1}$ is the sum of all irreducible characters $\chi_{\mu}$ such that $\mu$ is obtained from $\lambda$ by removing a removable box. By Frobenius reciprocity, the induction of $\chi_{\lambda}$ to $S_{n+1}$ is the sum of $\chi_{\mu}$ for $\mu$ obtained from $\lambda$ by adding an addable box.
@@ 2.2 General linear groups
The complex character theory of the general linear groups $\operatorname{GL}_{n}(q)$ uses many of the combinatorial concepts involved in symmetric groups such an hook lengths. One may find a description of some of this theory in [2] and [11].
We consider a subset of the ordinary characters, called unipotent characters. Although the general definition for arbitrary groups of Lie type is complicated, for $\operatorname{GL}_{n}(q)$ there is an easier description of these: they are the constituents of the permutation character on the Borel subgroup of upper triangular matrices.
These characters can be labelled by partitions of $n$, and their degrees are given by polynomials in $q$, called the generic degree of the character (or of the corresponding partition). If $\lambda=(\lambda_{1},\dots,\lambda_{r})$ is a partition of $n$, define Lusztig’s $a$-function by
@@
Lusztig’s $a$-function can be described as the smallest possible sum of entries in a semistandard tableau of shape $\lambda$ (with entries from $\{0,1,\dots\}$). The degree of the character $\chi_{\lambda}$ is (see [11, 3.2.6])
@@
Although two unipotent characters for $\operatorname{GL}_{n}(q)$ for a fixed $q$ could coincidentally have the same character degree, we are more interested in those partitions $\lambda$ and $\mu$ for which the generic degrees coincide, not just their value at a particular $q$. Notice that $\chi_{\lambda}(1)=\chi_{\mu}(1)$ if and only if $a(\lambda)=a(\mu)$ and the hook length multisets $H(\lambda)$ and $H(\mu)$ coincide. Thus this is a strictly stronger question than whether the symmetric group character degrees coincide. In general very little is known about the coincidence of generic degrees of unipotent characters. [^1] If you are wondering how this reconciles with the statement in the Introduction that it was relatively easy to find many characters of groups of Lie type of the same degree, the characters with the same degree found in [20] are not unipotent.
In geometric and combinatorial versions of Broué’s conjecture, there is a bijection between the irreducible unipotent characters in a given block $B$ of a group of Lie type and those of its Brauer correspondent $b$. This bijection can be found using a deformation of the algebra $b$, called a cyclotomic Hecke algebra. The parameters of the cyclotomic Hecke algebra are known, and to each character of the cyclotomic Hecke algebra there is a generic degree attached to it, calculated from the parameters, which is the generic degree of one of the unipotent characters in the block $B$. If all of the generic degrees of the characters in $B$ are distinct, then there is a canonical bijection between the characters of $B$ and of $b$, going via this cyclotomic Hecke algebra. Proposition 3.7, and particularly the data in Table 4.2, show that one cannot easily read off this bijection, and one actually has to calculate it, at least for some of the characters. [^2] To check if two unipotent characters lie in the same block module $\ell$, we first find the multiplicative order $d$ of $q$ modulo $\ell$, and then $\chi_{\lambda}$ and $\chi_{\mu}$ lie in the same $\ell$-block if and only if $\lambda$ and $\mu$ have the same $d$-core. If $d=1$ then this condition even becomes vacuous, but there will be other $d$ where the $d$-core is self-conjugate. Thus the results on unipotent character degree coincidences have impact on local-global conjectures in modular representation theory (see [5, 6] for more details).
@@ 3.1 Symmetric groups
For (ii), we find nine periodic clusters of period $9$, order $8$ and various sizes, one for each congruence modulo $9$.
@@
It is not known whether there are clusters of order $8$ of these sizes, but it is certainly likely that there are clusters of some of the larger sizes in the list above. There is no cluster of order $8$ and size up to $83$ by an exhaustive computer search, but computer programs take increasingly inconveniently long times to prove this. [^3] The case $n=82$ took around two weeks using a Magma program, just because of the sheer number of partitions involved. By $n=90$, the largest case known, it takes months, with no real option for parallelization. From an exhaustive search, there is a single cluster of order $8$ of sizes $84$ and $85$ (all other clusters have order $4$ or smaller), and no cluster of order greater than $4$ of sizes $86$ to $90$.
@@
Moving from the multisets having to coincide to coincidences of character degrees, we can prove much better bounds than those in Theorem 3.1. First, for $n\leq 129$ we can use Magma to explicitly calculate $m(n)$. [^4] The reason we stop at $n=129$ is technical. This is the point at which the author’s programs start to break down, because we start reaching what appear to be limits on the maximum size of a set in Magma, or something like the maximum number of bytes available to store a set. It was not processor or memory limits. One could go slightly higher by splitting the set of character degrees up and repeatedly running the program, but such manoeuvres will only eke out a few more cases, with one having to slice the set of character degrees up in ever finer ways. One might need to go to, perhaps, $n=200$ to see a better pattern in Figure 3.1, and that is simply infeasible. It is given in Table 4.1. Combining this with Theorem 3.1 proves almost all of the next result.
@@
For the bound for $m(n)\geq 4$, we can use Theorem 3.1 to prove the result for $n\geq 22$, and then Table 4.1 for the remaining values of $n$. For the second and third parts, Theorem 3.1 proves the result for $n\geq 206$, and from the remark afterwards, we know the result for $n\geq 182$. Table 4.1 gives us the result for $n\leq 129$. At this point one can either check each value in the range $130\leq n\leq 181$ individually that $m(n)\geq 8$ just by calculating all character degrees one at a time until at least eight of the same are found, or use the other known clusters of order $8$ to reduce the number of possible $n$ for which a brute-force search is needed. Either way, this proves the result.
@@
From Table 4.1 we see what the answer should be for larger bounds: for $m(n)\geq 10$ we need $n\geq 25$, for $m(n)\geq 12$ we need $n\geq 28$, and so on. However, this is conjectural because we do not have a version of Theorem 3.1 for these values, to eliminate all but a small number of $n$ that we can check individually. The proof in [3] gives no explicit bounds, and the version in [4] gives us explicit bounds of the form $m(n)\geq O(n^{1/7})$. These are obviously not useful for larger values of $m(n)$.
@@
However, we know that there are coincidences of character degrees, since $m(n)\to\infty$ as $n\to\infty$. Figure 3.1 gives the ratio $k(S_{n})/|\mathrm{cd}(S_{n})|$ for $1\leq n\leq 129$. It looks far from conclusive at this stage whether the average multiplicity has a limit (maybe 2.5? [^5] A suggestion from one of the referees is that it could possibly be related to the constant $\pi\sqrt{2/3}\sim 2.565$, since this quantity appears in the asymptotic formula for the number of partitions, $p(n)\sim\frac{e^{\pi\sqrt{2n/3}}}{4n\sqrt{3}}.$ ) or tends to infinity as $n$ increases. [^6] In [20], Moretó calculated this ratio for $n\leq 55$, but the general trend is not particularly obvious for those smaller values of $n$.
[…]
Graph of the number of conjugacy classes divided by the number of distinct character degrees for n up to 129. The graph is quite jagged, but increases overall but more slowly as n increases. For n above 100 the graph is slightly jagged and very slowly increasing.
@@
We can ask the question for arbitrary finite groups, a significant strengthening of the result by Moretó [20] and the author [3] on $|G|$ being bounded by a function of $m(G)$. In [20, p.245], Moretó asks this question specifically for soluble groups rather than all groups, but it looks as though this might be true, at least for all groups other than symmetric groups.
@@
There exists a function $f:\mathbb{Q}\to\mathbb{Q}$ such that, if $G$ is a finite group with no alternating composition factor, then $|G|\leq f(k(G)/|\mathrm{cd}(G)|)$.
The only reason that alternating groups are excluded from this is that Figure 3.1 does not compel the author to believe that the average multiplicity is unbounded for symmetric groups. We discuss for which groups this conjecture is known below, after the next conjecture.
@@
There exists a function $f:\mathbb{Q}\to\mathbb{Q}$ such that, if $G$ is a finite group then $|G|\leq f(k(G)/|\mathrm{cc}(G)|)$.
[…]
@@
We note that Conjectures 3.3 and 3.4 hold for $p$-groups and soluble groups with trivial Frattini subgroup [21, Lemmas 2.2 and 2.3], and Lie-type groups of bounded rank by [20, Lemma 6.3]. Conjecture 3.3 is true for simple groups of Lie type by [15, Corollary 1.6].
Having discussed the first column of the character table, one could ask the question about more than one column. A question of Alexander Ryba was whether there are irreducible characters $\lambda$ and $\mu$ of $S_{n}$ for $n\geq 2$ such that $\lambda(1)=\mu(1)$ and $\lambda(g)=\mu(g)\neq 0$ for any $g$ of cyclic type $2,1^{n-2}$ (so that $\lambda$ and $\mu$ are not conjugate)? The first such case is for $S_{12}$, where the characters associated to the partitions $(7,1,1,1,1,1)$ and $(4,4,4)$ have the same character degree and value on an involution. By $S_{30}$ there are such 70 pairs of partitions, and by $S_{40}$ there are 314. The first $n$ for which there are two characters with the same value on the conjugacy classes of $1$, $(1,2)$ and $(1,2,3,4)$ simultaneously is $35$.
@@
Let $X$ be a finite set of permutations and let $k\geq 1$ be an integer. For all sufficiently large $n$, there exist $k$ distinct irreducible characters of $S_{n}$ whose values on permutations $g\in S_{n}$ coincide for all $g$ of cycle type in $X$.
This conjecture contrasts with a recent result of Miller [19] which is that, in some sense, you don’t need many character values of $\chi^{\lambda}$ to determine $\lambda$. Why this does not contradict this conjecture is because in Miller’s algorithm you can choose the next class to examine the character value for based on the previous known character values, whereas in the conjecture you have to choose the classes in advance.
A result of Wildon [27, Proposition 2.1] states that if $\chi$ and $\psi$ are irreducible characters of $S_{n}$, and $\chi$ and $\psi$ agree on all $i$-cycles for $1\leq i\leq n$, then $\chi=\psi$. Thus we need only consider cycles in the previous conjecture, where the Murnaghan–Nakayama and Frobenius formulae are of a simpler form. Work around Kerov’s character polynomials (see [8, Section 1.2]), which specifically consider character values on cycles, might be of use when attacking this problem.
There is also an alternative formulation of the previous conjecture that seems more attractive. That $|S_{n}|$ is bounded in terms of $m(n)$ is the case $m=1$ of this next conjecture. It is unclear to the author whether even the case $n=2$ follows from the construction for the case $m=1$ from [3].
Conjecture 3.6.
Let $m,k\geq 1$. For all sufficiently large $n$, there are $k$ (distinct) irreducible characters of $S_{n}$ whose restrictions to $S_{m}$ coincide.